# Gradient elasticity in Swift-Hohenberg and phase-field crystal models

### v1

@@ 6 Conclusion
We discussed how gradient elasticity (GE) is encoded in Swift-Hohenberg (SH) and phase-field crystal (PFC) models. At leading order for small deformations, it consists of a first strain-gradient elasticity formulation, anisotropic for stripe phases and isotropic for crystalline phases. We derived analytical or semi-analytical formulations for GE characteristic lengths for one and multimode approximations of the periodic order parameter for different crystalline arrangements in a simplified setting. Moreover, we outlined the possibility of extending the parametrization via the design of the nonlocal terms in the free energy.
@@
The computed characteristic lengths are in the order of fractions of the lattice parameter, i.e., in the order of 1Å as observed for several materials. From a quantitative point of view, they are consistent with results reported in the literature, although they underestimate the ones commonly found for real materials with the same lattice symmetries [17, 74]. Moreover, in the minimal formulation of SH and PFC models, they can only be slightly varied, conforming to the known restrictions on the tuning of elastic constants in these approaches.
@@
However, numerical simulations which inherently encode a coarse-graining over microscopic lengths and do not consider any approximation, show that larger effective GE characteristic lengths emerge. They can be significantly varied with the model parameters. In particular, we focused on the elastic field at dislocations, which is of central interest for both GE and SH/PFC models. Interestingly, the solution of dislocation fields in both first and second strain-gradient elasticity can be fitted well to the results of APFC simulations. By analyzing the values of the emerging effective characteristic lengths, we found GE-2 to be a better model for describing the simulated stress field, thus assessing the underlying GE theory. Overall, we established that these theories are suitable for reproducing the elastic field of dislocations within the considered models, including regularization at the dislocation cores. On the other hand, we show that these GE theories naturally emerge from the minimal framework of the SH energy functional, which can be considered the simplest free-energy form minimized by a periodic, smooth order parameter [32].
@@
The effective GE characteristic lengths at dislocations vary with the model parameters similarly to the correlation length. This closely resembles variations observed for the size of defect cores in smooth theories for ordered systems [79]. Importantly, the variation in the quenching depth can be interpreted as a dependence on temperature. Thanks to the derivation from SH and PFC models, GE is directly linked to a framework for order-disorder phase transition. The dependence on quenching depth may constitute the input for other theories.
@@ Acknowledgements
We acknowledge interesting discussions with Ken R. Elder, Markus Lazar, and Axel Voigt. We also gratefully acknowledge support from the German Research Foundation under Grant No. SA4032/2 – Emmy Noether Programme – (MS) and SA4032/3 – FOR3013 – (LMB), and from the Center for Information Services and High-Performance Computing [Zentrum für Informationsdienste und Hochleistungsrechnen (ZIH)] at TU Dresden for computing time.
@@
The data that support the findings of this study are openly available at the following URL/DOI: $<$Information provided in the next version$>$.
@@ Appendix A Numerical Method
with $\eta_{n}$ the amplitudes to solve for, $\mathfrak{L}$ the linear operator in $\eta_{n}$ of Eq. (35), and $\mathfrak{N}$ the nonlinear part. We recall that the latter includes nonlinear terms in all the amplitudes $\eta_{n}$. In the Fourier pseudo-spectral method, we then solve for
@@
$\begin{split}\sigma_{ij}=\sum_{m}^{M}&\bigg{\{}\big{[}(\partial_{i}+{\rm i}(\mathbf{k}_{m})_{i})(\nabla^{2}+2{\rm i}\mathbf{k}_{m}\cdot\nabla)\eta_{m}\big{]}\big{[}(\partial_{j}-{\rm i}(\mathbf{k}_{m})_{j})\eta_{m}^{*}\big{]}\\ &-\big{[}(\nabla^{2}+2{\rm i}\mathbf{k}_{m}\cdot\nabla)\eta_{m}\big{]}\big{[}(\partial_{i}-{\rm i}(\mathbf{k}_{m})_{i})(\partial_{j}-{\rm i}(\mathbf{k}_{m})_{j})\eta_{m}^{*}+{\rm c.c.}\big{]}\bigg{\}},\end{split}$ (63)
@@ Appendix C Details of fitting procedures for characteristic lengths
The equations reported in B, upon centering the frame of reference at the dislocation core in $(x_{0},y_{0})$, illustrate functions $\sigma_{ij}(x,y,x_{0},y_{0},E,\nu,\xi_{1},\xi_{2})$ with $\xi_{i}=\ell_{i}$ for GE-1 and $\xi_{i}=\varrho_{i}$ for GE-2. These are here fitted (exploiting scipy.optimize [84]) on $\sigma_{ij}(x,y)$ in a squared region $15a_{0}\times 15a_{0}$ centered on a defect to determine $\xi_{i}$. Exploiting the elastic constants obtained in Table 1, and in agreement with [61, 40], we have that for the one-mode approximation of the triangular lattice the Lamé constants read $\mu=\lambda=3A\varphi_{1}^{2}$, which results in the Young modulus $E=5A\varphi_{1}^{2}/2$ and Poisson ratio $\nu=1/4$ (under the plane strain assumption). $\varphi$ can be computed analytically for one-mode approximations [47]. The value for the considered triangular lattice is $\varphi=(-C+\sqrt{C^{2}-15BD})/(15D)$. The dislocation in the APFC simulations moves slightly upon reaching equilibrium. In general, $x_{0}$ and $y_{0}$ are then to be considered as fitting parameters as well. The fit results for the fields illustrated in Fig. 3 are reported in Table 2. The fitted values of the characteristic lengths $\ell_{i}$ and $\varrho_{i}$ are indicated with the subscripts “I” and “II”.
@@
Table 2: Results of the fits illustrated in Fig. 3(a). Values $(\ell_{\rm I},\ell_{\rm II},x_{0}^{\text{GE-1}},y_{0}^{\text{GE-1}})$ are obtained by fitting Eqs (64)–(66) with characteristic lengths and positions as fitting parameters. The values $\ell_{\rm I}=\ell_{\rm II}$ are obtained from fitting the same equation but enforcing this constraint. Values $(\varrho_{\rm I},\varrho_{\rm II},x_{0}^{\text{GE-2}},y_{0}^{\text{GE-2}})$ are obtained by fitting Eqs (67)–(69), also in this case with characteristic lengths and positions as fitting parameters. All quantities are expressed in $a_{0}$ units (fraction of the lattice parameter). For all these values, the standard deviation does not exceed $2\times 10^{-4}$.
The values of fitted parameters are overall consistent across different stress components, with just GE-1 characteristic lengths obtained by fitting $\sigma_{xy}$ deviating from other components, as discussed in the main text. Note that $x_{0}$ differs from 00 by less than $5\%$ of the lattice parameter. The shift along the $x$ direction could then be considered negligible. Instead, we obtain a shift along $y$ of $\sim 20\%$ of the lattice parameter. Note that the dislocation core location $(x_{0},y_{0})$ is quantitatively consistent also when looking at the fitted values on both GE-1 and GE-2 stress field components, where indeed $x_{0}$ and $y_{0}$ have the same meaning. In the plots reported in the main text, the stress fields from GE solutions for dislocations are centered in the $(x_{0},y_{0})$ values obtained from fitting the corresponding stress component. We remark that in GE-2, the two length scales $\varrho_{1,2}$ are fully equivalent as it follows from Eq. (23), so their values could be exchanged. Also, in some parameter ranges, e.g., for small $A$ values, the best fit is given by $\varrho_{1}\sim\varrho_{2}$ while in other cases a significant difference, as in the Table above, is obtained. The physically meaningful lengths entering the theory are, however, $\omega(\varrho_{1},\varrho_{2})$ and $\gamma(\varrho_{1},\varrho_{2})$ as defined in Sect. 2.3 and these are the quantities illustrated in Fig. 4. In this figure, we also report the correlation length $W$ for triangular lattice when varying A and B. For this lattice symmetry in the one-mode approximation, an analytic solution connecting $W$ to the model parameter exists [78]
@@ Appendix D On higher order GE terms from nonconstant amplitudes
The main text discusses evidence of second-strain gradient elasticity contributions in numerical simulations. In particular, the stress field at the dislocation core and the detailed analysis by comparison with analytic solutions within GE-1 and GE-2 show that the latter better describes the simulation results.
In Sect. 3, we show that under the assumption of constant magnitude of the amplitude ($|\nabla\varphi_{n}|=0$) typically considered for small deformations, GE is encoded in the SH energy functional via first strain-gradient terms only. This assumption, however, does not strictly hold true in simulations [45, 80, 81]. While refraining from a complete derivation of the theory, we report here a simplified argument to show how including the amplitude gradients leads to second strain-gradient terms in the elastic energy.
@@
Therefore, $\varphi^{\prime\prime}(x)$ and, by extension, the elastic energy density, contain a contribution from $u^{\prime\prime\prime}(x)$, i.e. the second derivative of the strain field, which is the higher-order term entering GE-2 but absent in GE-1. As such, GE-2 effects may be ascribed to gradients in the amplitude, while their presence is anyhow corroborated by numerical results and, importantly, a detailed comparison with specific analytical solutions.

### v2

@@ 6 Conclusions
We discussed how gradient elasticity (GE) is encoded in Swift-Hohenberg (SH) and phase-field crystal (PFC) models. The leading order for small deformations consists of a first strain-gradient elasticity formulation, anisotropic for stripe phases and isotropic for crystalline phases. We derived analytical or semi-analytical formulations for GE characteristic lengths for one and multimode expansions of the periodic order parameter for different crystalline arrangements in an approximated setting.
@@
These characteristic lengths are in the order of fractions of the lattice parameter ($\sim$1Å) as observed for several materials. From a quantitative point of view, they are consistent with results reported in the literature. However, they underestimate the ones commonly found for real materials with the same lattice symmetries [19, 78]. When considering the minimal formulation of the SH energy functional, they can only be slightly varied, conforming to the known restrictions on tuning elastic constants in these approaches compared to fully atomistic methods. However, we outlined the possibility of extending the parametrization via the design of the nonlocal terms in the free energy. Although these results outline the leading-order GE effects, they are obtained via simplifying assumptions, namely neglecting nonlinear terms and assuming constant amplitude moduli everywhere.
@@
Numerical simulations that allow the inspection of GE effects without approximations show that larger effective GE characteristic lengths emerge. In addition, they can be significantly varied with the model parameters. We focused on the elastic field at dislocations, which is of central interest for both GE and SH/PFC theories. Interestingly, the analytical solutions of stress fields at edge dislocations in both first and second strain-gradient elasticity closely resemble the results of APFC simulations. By analyzing the values of the emerging effective characteristic lengths, we found GE-2 to be a better model for describing the simulated stress field, thus assessing the underlying GE theory. Overall, we showed that these GE theories, particularly GE-2, naturally emerge from the minimal framework of the SH energy functional, which can be considered the simplest free-energy form minimized by a periodic, smooth order parameter [35].
@@
The effective GE characteristic lengths at dislocations vary with the model parameters similarly to the phase correlation length. This closely resembles variations observed for the size of defect cores in smooth theories for ordered systems [85]. Importantly, the variation in the quenching depth can be interpreted as a dependence on temperature, which may constitute the input for other theories and establishes a direct link between GE and the PFC/SH framework for order-disorder phase transition.
@@ Acknowledgements
We acknowledge interesting discussions with Ken R. Elder, Markus Lazar, and Axel Voigt. We also gratefully acknowledge support from the German Research Foundation under Grant No. SA4032/2 – Emmy Noether Programme – (MS) and SA4032/3 – FOR3013 – (LBM), and from the Center for Information Services and High-Performance Computing [Zentrum für Informationsdienste und Hochleistungsrechnen (ZIH)] at TU Dresden for computing time.
@@
The data that support the findings of this study are openly available at the following URL/DOI: 10.5281/zenodo.10351669
@@ Appendix A Numerical method
with $\eta_{n}$ the amplitudes to solve for, $\mathfrak{L}$ and $\mathfrak{N}$ the linear operator and polynomial terms in $\eta_{n}$ of Eq. (32), respectively. We recall that $\mathfrak{N}$ includes nonlinear terms in all the amplitudes $\eta_{n}$. In the Fourier pseudo-spectral method, we then solve for
@@
$\begin{split}\sigma_{ij}=\sum_{n=1}^{N}&\bigg{\{}\big{[}(\partial_{i}+{\rm i}(\mathbf{k}_{n})_{i})(\nabla^{2}+2{\rm i}\mathbf{k}_{n}\cdot\nabla)\eta_{n}\big{]}\big{[}(\partial_{j}-{\rm i}(\mathbf{k}_{n})_{j})\eta_{n}^{*}\big{]}\\ &-\big{[}(\nabla^{2}+2{\rm i}\mathbf{k}_{n}\cdot\nabla)\eta_{n}\big{]}\big{[}(\partial_{i}-{\rm i}(\mathbf{k}_{n})_{i})(\partial_{j}-{\rm i}(\mathbf{k}_{n})_{j})\eta_{n}^{*}+{\rm c.c.}\big{]}\bigg{\}}.\end{split}$ (61)
@@ Appendix C Details of fitting procedures for characteristic lengths
The equations reported in B, upon centering the frame of reference at the dislocation core in $(x_{0},y_{0})$, illustrate functions $\sigma_{ij}(x,y,x_{0},y_{0},E,\nu,\xi_{1},\xi_{2})$ with $\xi_{i}=\ell_{i}$ for GE-1 and $\xi_{i}=\varrho_{i}$ for GE-2. They are fitted (exploiting scipy.optimize [88]) on $\sigma_{ij}(x,y)$ in a squared region $15a_{0}\times 15a_{0}$ centered on a defect to determine $\xi_{i}$. Exploiting the elastic constants obtained in Table 1, and in agreement with [61, 42], we have that for the one-mode approximation of the triangular lattice the Lamé constants read $\mu=\lambda=3A\varphi_{1}^{2}$, which results in the Young modulus $E=5A\varphi_{1}^{2}/2$ and Poisson ratio $\nu=1/4$ (under the plane strain assumption). $\varphi$ can be computed analytically for one-mode approximations [49]. The value for the considered triangular lattice is $\varphi=(-C+\sqrt{C^{2}-15BD})/(15D)$. The dislocation in the APFC simulations moves slightly upon reaching equilibrium. In general, $x_{0}$ and $y_{0}$ must be considered as fitting parameters as well. The fit results for the fields illustrated in Fig. 3 are reported in Table 2. The fitted values of the characteristic lengths $\ell_{i}$ and $\varrho_{i}$ are indicated with the subscripts “I” and “II”.
@@
Table 2: Results of the fits illustrated in Fig. 3(a). Values $(\ell_{\rm I},\ell_{\rm II},x_{0}^{\text{GE-1}},y_{0}^{\text{GE-1}})$ are obtained by fitting Eqs (62)–(64) with characteristic lengths and positions as fitting parameters. The values $\ell_{\rm I}=\ell_{\rm II}$ are obtained from fitting the same equation but enforcing equal characteristic lengths. Values $(\varrho_{\rm I},\varrho_{\rm II},x_{0}^{\text{GE-2}},y_{0}^{\text{GE-2}})$ are obtained by fitting Eqs (65)–(67), also in this case with characteristic lengths and positions as fitting parameters. All quantities are expressed in $a_{0}$ units (fraction of the lattice parameter). For all these values, the standard deviation does not exceed $2\times 10^{-4}$.
The values of fitted parameters are overall consistent across different stress components, with just GE-1 characteristic lengths obtained by fitting $\sigma_{xy}$ deviating from other components, as discussed in the main text. Note that $x_{0}$ differs from 00 by less than $5\%$ of the lattice parameter. The shift along the $x$ direction could then be considered negligible. Instead, we obtain a shift along $y$ of $\sim 20\%$ of the lattice parameter. Note that the dislocation core location $(x_{0},y_{0})$ is quantitatively consistent across the fitted values on both GE-1 and GE-2 stress field components, where indeed $x_{0}$ and $y_{0}$ have the same meaning. In the plots reported in the main text, the stress field from GE solutions for dislocations are centered in the $(x_{0},y_{0})$ values obtained from fitting the corresponding stress component. We remark that in GE-2, the two length scales $\varrho_{1,2}$ are fully equivalent as it follows from Eq. (23), so their values could be exchanged. Also, in some parameter ranges, e.g., for small $A$ values, the best fit is given by $\varrho_{1}\sim\varrho_{2}$ while in other cases a significant difference, as in Table 2 above, is obtained. The physically meaningful lengths entering the theory are, however, $\omega(\varrho_{1},\varrho_{2})$ and $\gamma(\varrho_{1},\varrho_{2})$ as defined in Sect. 2.3 and these are the quantities illustrated in Fig. 4. This figure also reports the correlation length $W$ for triangular lattice when varying $A$ and $B$. For this lattice symmetry in the one-mode approximation, an analytic solution connecting $W$ to the model parameter exists [84]
@@ Appendix D On higher order GE terms from nonconstant amplitudes
In Sect. 4, we show that under the assumption of constant magnitude of the amplitudes ($|\nabla\varphi_{n}|=0$) typically considered for small deformations, GE is encoded in the SH energy functional via first strain-gradient terms only. This assumption, however, does not strictly hold in numerical simulations [47, 89, 90]. While refraining from a complete derivation of the theory, we report here a simplified argument to show how including the amplitude gradients leads to second strain-gradient terms in the elastic energy.
@@
Therefore, $\varphi^{\prime\prime}(x)$ and, by extension, the elastic energy density, contain a contribution from $u^{\prime\prime\prime}(x)$, i.e. the second derivative of the strain field, which is the higher-order term entering GE-2 but absent in GE-1. As such, GE-2 effects may be ascribed to gradients in the amplitude, while their presence is anyhow corroborated by numerical results and, importantly, a detailed comparison with specific analytical solutions as discussed in Sect. 5.